An interest rate of 17,5% per year, compounded quarterly, is equivalent to a continuous compounding rate of

[1] 17,185%.

[2] 17,500%.

[3] 17,128%.

[4] 19,125%.

[5] 17,888%


1
Expert's answer
2021-05-28T05:13:18-0400

Continuous compound interest is based on the fact that a division of the rate by n as n approaches infinity is


(1+r/n)n = er (When r=1 this comes to e itself.)


This gives rise to the continuous compound interest formula


A=Pert,

where t is time, r is rate, P is the principal amount and A the amounts after time t. If we take t to be the number of quarters and A/P-1 to represent the rate of interest earned after t quarters, then




ert - 1 = e(0.175/4*4) - 1


=1.191246-1


= 0.191246


=19.125%




Therefore, answer is (4)